# Learning Unmasking Policies for Diffusion Language Models

### v2

@@ 4.3 Transferability of RL Sampling Policies
We next turn to the question of transferability. A notable advantage of heuristic approaches like Fast-dLLM is that they can be applied post-hoc to any model or dataset. [^4] Though heuristics might still require manual hyperparameter tuning as exemplified by different optimal threshold across datasets for Fast-dLLM (e.g., comparing Figure 3(a) vs Figure 4(b)). This naturally raises the question: to what extent can RL policies trained on a specific model and dataset be reused on other models/datasets?
@@
We begin by investigating to what extent the policies found via RL transfer across models; note that such transfers are possible because our policies rely solely on token confidences and are therefore agnostic to token embeddings, which would not transfer from one model to another. We thus reuse the policies trained on the LLaDA-8B-Instruct model (under the semi-AR setting; cf. Section˜4.1) and evaluate them on Dream-7B-Instruct; the results are shown in Figure˜4(a) (see Figure˜10 for MATH). Encouragingly, most of the LLaDA-trained policies nearly match the performance of Fast-dLLM when evaluated on Dream, and perform very similarly to those trained on Dream directly (Figure 7). The one exception is the $\alpha=10$ policy ( ), which collapses to Fast-dLLM performance and fails to retain the good low-NFE performance observed with LLaDA (cf. Figure˜3(a))—suggesting that the extreme steepness of the reward curve (due to high $\alpha$) causes this policy to overfit to model-specific patterns in LLaDA’s confidence levels.
Domain transfer: Mathematical reasoning to code.
Next, we examine how well our policies transfer across different data domains. We again reuse the policies from Section˜4.1, which were trained on a mixture of mathematical data (GSM8K/MATH), but this time evaluate them on the coding tasks of HumanEval (chen2021evaluating) and MBPP (austin2021program). As shown in Figure˜4(b) and Figure˜4(c) ( ), we find that these policies fail to fully transfer to the coding domains: their performance more closely aligns with the high-confidence baseline rather than with Fast-dLLM. To investigate whether this drop in performance is due to the lack of domain-relevant training data, we train a new policy on the coding dataset KodCode-RL-10K (xu2025kodcode) ( ). These coding-specific policies narrow the gap to the Fast-dLLM baseline on HumanEval and MBPP, underscoring the importance of using a diverse data mixture to support generalization across domains. We leave to future work the training of sampling policies on mixtures that combine datasets from different domains (e.g., mathematics and coding).
@@
(d) Sequence length transfer: $256\to 512$ on GSM8k.
@@
Lastly, we examine whether our policies transfer across different sequence lengths $L$ (not to be confused with the block length $BL$ used in semi-AR generation). Such transfer is possible since we instantiate $f_{\theta}$ with a transformer trained with rotary positional embeddings (su2021roformer), rather than, for example, an MLP or other fixed-length architectures. We thus take the policies from Section˜4.2 and evaluate them at a 2x longer sequence length, $L=512$. Results are shown in Figure˜4(d) (see Figure 11 for MATH results). While the baselines degrade further as the sequence length increases, the learned policies yield similar performance to before, suggesting that RL policies can transfer effectively across generation lengths without retraining.
@@ 4.4 Exploring the Design Space of $\pi_{\phi}$
Lastly, we conduct a series of experiments to elucidate the impact of various design choices behind our proposed RL sampling method.
Additive vs. multiplicative rewards
@@
We begin by ablating the structure of the reward function, comparing our proposed multiplicative combination of correctness and computational terms (cf. Equation˜3.1) with an additive alternative: $r({\bm{y}},{\bm{y}}_{\hat{T}})-\alpha\big(\frac{T-\hat{T}}{T}\big)$. As shown in Figure˜5(a), while both exhibit increasing reward as training progresses, we observe that the additive formulation is much more prone to ‘reward hacking’, where it collapses to a very-fast-but-very-wrong policy. This is illustrated in Figure˜5(b): training with the additive reward results in a policy that unmasks everything at once, leading to the number of sampling steps to collapse to the minimum possible for all samples (i.e., $8$ steps for training with $BL=32$ and $L=256$). As discussed in Section˜3.3, we attribute this issue to the fact that incorrect but fast samples may still receive a positive advantage under the additive reward. In contrast, we find that the multiplicative reward effectively mitigates this issue by assigning a positive advantage only if the generation is correct, resulting in more stable and predictable training behavior.
Policy likelihood
@@
Next, we examine our choice of the policy likelihood parameterization. Recall from Section˜3.2 that we model the unmasking probability for each position using a Bernoulli distribution. This has the advantage of admitting very efficient inference and likelihood calculations, but relies on the network $f_{\phi}$ to implicitly embed relationships between different tokens in the scores ${\bm{b}}_{t}$, and runs the risk of producing ${\mathcal{U}}_{t}^{\pi}=\emptyset$ in case ${\bm{u}}_{t}=\mathbf{0}$. In this experiment, we therefore investigate a more involved sampling procedure, which we call dynamic Plackett-Luce sampling (DPLS; detailed description in Appendix C) which is guaranteed to unmask at least one position in each step and which non-linearly combines the scores ${\bm{b}}_{t}$ through a softmax, while still allowing for variable-size unmasking sets ${\mathcal{U}}_{t}^{\pi}$. Since this affects not only the sampling strategy but also the likelihood computation, we retrain the policies from Section˜4.1; the resulting downstream accuracy on GSM8k is then showed in 5(c). We observe that both methods achieve very similar performance, aligning closely with the Fast-dLLM frontier, which might provide additional support for the hypothesis that this frontier is optimal in the semi-AR setting. Furthermore, DPLS policies appear to show slightly better controllability via $\alpha$ parameter (as indicated by a larger spread of policies trained with varying $\alpha$).
Confidence-based policy input
Finally, we revisit our design choice of relying solely on the maximum token confidence values $c_{t}^{k}$ as input to the sampling policy (cf. Section˜3.2) to explore whether other measures of uncertainty in the base model’s predictions could yield better results. We first train and evaluate a set of policies which do not take only the highest confidence per position $c_{t}^{k}$ as input, but rather the top 50 highest values, thus giving the model more detailed information about the token predictive distributions $p_{t}^{k}$ and potentially allowing it to design its own confidences measures for effective sampling. The results are presented in Figure˜5(d). Somewhat surprisingly, using only the maximum confidence per position ( ) performs slightly better than using the top-50 confidences ( ), which suggests that alternative uncertainty measures are unlikely to yield performance gains over the simple confidences $c_{t}^{k}$.
@@
(b) Mean number of NFEs when training LLaDA with additive vs. multiplicative reward (both $\alpha=1.0$).
(c) Bernoulli vs DPLS sampling on GSM8k. Both for LLaDA on GSM8k, $\tau_{\pi}=0.5$.
(d) Bernoulli policy with top-1 versus with top-50 confidences. Both for LLaDA on GSM8k, $\tau_{\pi}=0.5$.
@@
Additionally, we consider parametrizing the policy as an additional classification head on top of LLaDA’s final hidden state ${\bm{h}}_{t}^{k}\in\mathbb{R}^{H}$. While this results in a significantly larger policy ($300M$ parameters, an increase of $1,000\times$), it offers the potential advantage of incorporating token-level semantic information. However, in practice we observe that hidden state-based policies perform worse (see in Figure˜5(d)) and exhibit less stable training dynamics than the confidence-based policies. These results suggest that the unembedding matrix $W\in\mathbb{R}^{H\times V}$, which maps hidden states to token logits, plays a vital role in enabling effective policy decisions via confidence signals. [^5] Similar findings have been reported in the early-exit literature (schuster2022confident), where confidence-based stopping criteria have been shown to significantly outperform those relying on hidden state representations.
@@ 5 Related work
Sampling in diffusion LLMs (dLLMs) (llada2025; ye2025dream) has recently attracted significant attention, with much of the work in this area proposing heuristic approaches to improve the decoding process in a training-free manner. Throughout this paper, we focus on Fast-dLLM (fastdllm2025) as a representative heuristic method, as it popularized confidence-thresholded sampling in dLLMs and demonstrated its crucial role in enabling faster inference in dLLMs compared to autoregressive models. Many recently proposed heuristics can be viewed as either reinterpretations or extensions of such confidence thresholding strategies (ben2025accelerated; wei2025accelerating; hong2025wide; li2025diffusion; yu2025dimple). Other notable heuristic approaches explore incorporating spatial (huang2025pc) or temporal information (wang2025time), alternative confidence measures (kim2025train), or more explicit modeling of token dependencies (azangulov2025parallel). In this work, we aim to complement ongoing research on sampling heuristics by investigating whether effective sampling strategies can be learned directly via reinforcement learning. Note that beyond improving the efficiency of unmasking in dLLMs, heuristics have also been proposed for remasking (hong2025wide; dong2025saber) and for dynamically adjusting the generation length (li2025beyond), which we do not target in this work.
@@
Early work on post-training diffusion LLMs via reinforcement learning includes d1 (zhao2025d1), which introduces a variant of GRPO tailored to dLLMs, and DiffuCoder (gong2025diffucoder), which focuses on enhancing the coding abilities of LLaDA-style models using RL. Most recent extensions aim to improve the quality of policy gradient estimators (tang2025wd1; wang2025spg; lin2025boundary; rojas2025improving; zhu2025enhancing; wang2025revolutionizing; zhan2025principled), and have demonstrated promising results in further enhancing the reasoning capabilities of dLLMs. The key distinction from our approach is that these methods use a fixed sampling strategy (e.g., high-confidence sampling), and the policy corresponds to the dLLM itself (or a LoRA-augmented version for efficiency). Closest to our work are DCOLT (huang2025reinforcing), which trains a separate unmasking module in addition to updating the base model via RL, and DiFFPO (zhao2025diffpo), which jointly learns an unmasking confidence threshold while updating the base model through RL. Differently to both these works, our primary goal is not in improving the reasoning abilities of the underlying dLLM; hence we keep the underlying dLLM fixed and focus solely on training a standalone policy with the aim of learning fast sampling while preserving the performance of the base model. Finally, in concurrent work, Seed Diffusion (song2025seed) proposes computation-aware RL as one of the key ingredients for achieving competitive efficiency among closed-source coding dLLMs.
@@
Besides the aforementioned heuristic-based sampling innovations, other efforts to improve inference efficiency in dLLMs (ma2025dinfer) include KV caching (jiang2025d; fastdllm2025), (variants of) speculative decoding (israel2025accelerating; campbell2025self; guo2025reviving), training separate decoder modules during pretraining (liu2024think; arriola2025encoder), and diffusion forcing (wang2025diffusion), among others. Concurrent work explores distilling faster generation patterns directly into the base model (chen2025dparallel), or training a (small) separate unmasking network on top of the pretrained dLLM (bansal2025enabling; bao2025learning). Crucially, unlike our approach, where this training is done via RL, these works train unmasking modules by distilling generation trajectories from the base model. As a result, we expect that such approaches may still inherit limitations observed in dLLM sampling, such as difficulties in the non semi-AR regime (cf. Section˜2.2).
@@
Since our sampling policies result in a variable number of sampling steps ($T-\hat{T}$) per input, our work connects to the broader literature on adaptive computation (graves2016adaptive; teerapittayanon2016branchynet). Notable examples of using RL to learn input-dependent compute policies include conditional computation with stochastic gating (bengio2015conditional), dynamic block skipping in residual networks (wang2018skipnet), and RL-based early-exit decisions for long chain-of-thought reasoning (dai2025s). To the best of our knowledge, our work is the first to explore learning adaptive policies using RL for the task of sampling in dLLMs. The closest related concurrent work is DiFFPO (zhao2025diffpo); however, unlike our approach–where the policy is learned end-to-end–DiFFPO learns to predict adaptive thresholds, which are then used in the same fashion as the fixed thresholds employed by Fast-dLLM (fastdllm2025).
@@ 6 Conclusion
We introduced a reinforcement learning approach for learning unmasking strategies in diffusion LLMs. Our experiments demonstrated that the learned policies can match and sometimes exceed the performance of recently proposed sampling heuristics, paving the way for the automated discovery of scalable and robust sampling mechanisms.
Limitations and future work
While we have shown that training sampling policies on LLaDA (llada2025) allows us to match the performance of heuristics such as Fast-dLLM (fastdllm2025), a small performance gap remains when policies are trained on Dream (ye2025dream). Understanding which characteristics of Dream (e.g., it being initialized from an AR model) contribute to the diminished performance of RL policies is an important open question. One possible direction for future work is thus to train policies on a mixture of models, which might help bridge this gap, or to explore other lightweight ways of incorporating semantic information (e.g., by pairing confidences with the (un)embedding vectors of the corresponding tokens). Another limitation of our approach is that it requires training a separate policy for each value of $\alpha$ (cf. Equation˜3.1), and that this hyper-parameter sometimes does not yield as much fine-grained control over the policy’s behavior as desired (c.f. Section˜4.1). Another interesting avenue for future work is thus to investigate whether the accuracy-speed tradeoff could be controlled through some alternative mechanism, instead of explicitly having to set the strength of the computational penalty during training. We also observe that the optimal policy temperature $\tau_{\pi}$ varies across generation settings (cf. Figure˜3), suggesting the need to explore whether $\tau_{\pi}$ can be learned jointly with the policy, or whether alternative likelihood parameterizations (such as DPLS, see Appendix C) offer greater robustness to temperature selection. Beyond overcoming these limitations, promising future extensions include: (i) expanding the training data mixture to incorporate samples from multiple domains; (ii) extending our sampling policies to support not only unmasking but also remasking (wang2025remasking); and (iii) moving beyond the text domain to learn samplers for multimodal discrete diffusion models (swerdlow2025unified).
@@
  • In Appendix A, we present additional plots to supplement the experiments from the main paper:
    • In A.1, we replicate Figure˜1 across models (LLaDA, Dream) and datasets (GSM8k, MATH).
    • In A.2, we replicate Figure˜3 for policies trained on Dream.
    • In A.3, we show results for the same setting as 3(a) and 3(c) but with a denser $\alpha$-grid.
    • In A.4, we plot the impact of varying the policy temperature $\tau_{\pi}$.
    • In A.5, we provide model transfer results for the MATH dataset.
    • In A.6, we provide generation length (L) transfer results for the MATH dataset.
  • In Appendix B, we provide implementation and hyperparameters details.
  • In Appendix C, we describe dynamic Plackett-Luce sampling as an alternative to Bernoulli sampling.
  • In Appendix D, we detail our expert steering approach.
  • In Appendix E we discuss an extended background discussion on discrete diffusion models, to benefit readers who are not familiar with this paradigm.
Appendix A Additional Results
A.1 Figure 1 replicated for {LLaDA-8B-Instruct, Dream-7b-Instruct} $\times$ {GSM8k, MATH-500}

### v3

@@ 4.3 Transferability of RL Sampling Policies
We next turn to the question of transferability. A notable advantage of heuristic approaches is that they can be applied post-hoc to any model or dataset. [^1] Though heuristics might still require manual hyperparameter tuning, as exemplified by different optimal threshold across datasets for Fast-dLLM (e.g., comparing Figure 4(a) vs Figure 7(b)). This naturally raises the question: to what extent can RL policies trained on a specific model and dataset be reused in other settings?
@@
We begin by investigating the transferability of RL policies across models; note that such transfers are possible because our policies rely solely on token confidences and are therefore agnostic to token embeddings, which would not transfer from one model to another. We thus reuse the policies trained on LLaDA (cf. Section˜4.1) and evaluate them on Dream; the results are shown in Figure˜7(a) (see Figure˜15 for MATH). Encouragingly, most of the LLaDA-trained policies nearly match the performance of Fast-dLLM when evaluated on Dream, and perform very similarly to those trained on Dream directly (Figure 11). The one exception is the $\alpha=10$ policy ( ), which collapses to Fast-dLLM performance and fails to retain the good low-NFE performance observed with LLaDA (cf. Figure˜4(a))—suggesting that the extreme steepness of the reward curve (due to high $\alpha$) causes this policy to overfit to model-specific patterns in LLaDA’s confidence levels.
@@
(d) Sequence length transfer: $256\to 512$ (GSM8k).
@@
Domain transfer: Mathematical reasoning to code.
Next, we examine how well our policies transfer across different domains. We again reuse the policies from Section˜4.1, which were trained on a mixture of mathematical data (GSM8k/MATH), but this time evaluate them on the coding tasks of HumanEval (chen2021evaluating) and MBPP (austin2021program). As shown in Figure˜7(b) and Figure˜7(c) ( ), we find that these policies fail to fully transfer to the coding domains, especially on HumanEval. To investigate whether this drop in performance is due to the lack of domain-relevant training data, we train a new policy on the coding dataset KodCode-RL-10K (xu2025kodcode) ( ). These coding policies narrow the gap to the Fast-dLLM baseline on HumanEval and improve low-NFE performance on MBPP, underscoring the importance of using a diverse data mixture to support generalization across domains.
@@
Lastly, we examine whether our policies transfer across different sequence lengths $L$ (not to be confused with the block length $BL$ in semi-AR). Such transfer is possible since we instantiate $f_{\theta}$ with a transformer, rather than, for example, an MLP (which requires fixed-length inputs). We thus take the policies from Section˜4.2 and evaluate them at a 2x longer sequence length ($L=512$). Results are shown in Figure˜7(d) (see Figure 16 for MATH). While the baselines degrade further as the sequence length increases, the learned policies yield similar performance to before, suggesting that RL policies can transfer effectively across generation lengths without retraining.
@@ 4.4 Exploring the Design Space of $\pi_{\phi}$
Additive vs. multiplicative rewards.
@@
We begin by ablating the structure of the reward function, comparing our proposed multiplicative combination of correctness and computational terms (cf. Equation˜3.1) with an additive alternative: $r({\bm{y}},{\bm{y}}_{\hat{T}})-\alpha\big(\frac{T-\hat{T}}{T}\big)$. As shown in Figure˜8(a), while both exhibit increasing reward as training progresses, we observe that the additive formulation is much more prone to ‘reward hacking’, where it collapses to a very-fast-but-often-wrong policy. This is illustrated in Figure˜8(b): training with the additive reward results in a policy that unmasks everything at once, leading the number of sampling steps to collapse to the minimum possible for all inputs. As discussed in Section˜3.3, we attribute this issue to the fact that incorrect but fast samples may still receive a positive advantage under the additive reward. In contrast, we find that the multiplicative reward effectively mitigates this issue by assigning a positive advantage only if the generation is correct, resulting in more stable and predictable training behavior.
Policy likelihood.
@@
Recall from Section˜3.2 that we model the unmasking probability for each position using a Bernoulli distribution. This has the advantage of admitting very efficient inference and likelihood calculations, but relies on the network $f_{\phi}$ to implicitly embed relationships between different tokens in the scores ${\bm{b}}_{t}$, and runs the risk of producing ${\mathcal{U}}_{t}^{\pi}=\emptyset$ in case ${\bm{u}}_{t}=\mathbf{0}$. Here, we therefore investigate a more involved sampling procedure, which we call dynamic Plackett-Luce sampling (DPLS; detailed description in Appendix E) which is guaranteed to unmask at least one position in each step and which combines the scores ${\bm{b}}_{t}$ through a softmax. We retrain the policies from Section˜4.1 using DPLS; the resulting downstream accuracy on GSM8k is then shown in 8(c). We observe that both methods achieve very similar performance, aligning closely with the Fast-dLLM, which might provide additional support for the hypothesis that this frontier is optimal in the semi-AR setting. Furthermore, DPLS policies appear to show slightly better controllability via $\alpha$ (as indicated by a larger spread of policies trained with varying $\alpha$).
@@
(b) Mean NFEs when training LLaDA with additive vs. multiplicative reward (both $\alpha=1.0$).
(c) Bernoulli vs DPLS sampling. Both for LLaDA on GSM8k.
(d) Bernoulli policies with varying inputs. All for LLaDA on GSM8k.
@@
Confidence-based policy input.
Finally, we revisit our design choice of relying solely on the maximum token confidence values $c_{t}^{k}$ as input to the sampling policy (cf. Section˜3.2). We first train and evaluate a set of policies which do not take only the highest confidence per position $c_{t}^{k}$ as input, but rather the top 50 highest values, thus giving the model more detailed information about the token predictive distributions $p_{t}^{k}$ and potentially allowing it to design its own confidences measures for effective sampling. The results are presented in Figure˜8(d). Somewhat surprisingly, using only the maximum confidence per position ( ) performs slightly better than using the top-50 confidences ( ), which suggests that alternative uncertainty measures are unlikely to yield performance gains over the simple confidences $c_{t}^{k}$.
Additionally, we explore parametrizing the policy as an additional classification head on top of LLaDA’s final hidden state ${\bm{h}}_{t}^{k}\in\mathbb{R}^{H}$. While this results in a significantly larger policy ($300M$ parameters), it offers the potential advantage of incorporating token-level semantic information. However, in practice we observe that hidden state-based policies perform worse (see in Figure˜8(d)) and exhibit less stable training dynamics than the confidence-based policies. These results suggest that the unembedding matrix $W\in\mathbb{R}^{H\times V}$, which maps hidden states to token logits, plays a vital role in enabling effective policy decisions via confidence signals. [^2] Similar findings have been reported in the early-exit literature (schuster2022confident), where confidence-based stopping criteria have been shown to significantly outperform those relying on hidden state representations.
@@ 5 Related work
Sampling in diffusion LLMs (dLLMs) (llada2025; ye2025dream) has recently attracted significant attention, with much of the work in this area proposing heuristic approaches to improve the decoding process in a training-free manner. Throughout this paper, we focus on Fast-dLLM (fastdllm2025) as a representative heuristic method, as it popularized confidence-thresholded sampling in dLLMs and demonstrated its crucial role in enabling faster inference in dLLMs compared to autoregressive models. Many recently proposed heuristics can be viewed as either reinterpretations or extensions of such confidence thresholding strategies (ben2025accelerated; wei2025accelerating; hong2025wide; li2025diffusion; yu2025dimple; kim2025klass; shen2025improving). Other notable heuristic approaches explore incorporating spatial (huang2025pc) or temporal information (wang2025time), alternative confidence measures (kim2025train), or more explicit modeling of token dependencies (azangulov2025parallel). In this work, we aim to complement ongoing research on sampling heuristics by investigating whether effective sampling strategies can be learned directly via reinforcement learning. Note that beyond improving the efficiency of unmasking in dLLMs, heuristics have also been proposed for remasking (hong2025wide; dong2025saber) and for dynamically adjusting the generation length (li2025beyond), which we do not target in this work.
@@
Early work on post-training diffusion LLMs via reinforcement learning includes d1 (zhao2025d1), which introduces a variant of GRPO tailored to dLLMs, and DiffuCoder (gong2025diffucoder), which focuses on enhancing the coding abilities of LLaDA-style models using RL. Most recent extensions aim to improve the quality of policy gradient estimators (tang2025wd1; wang2025spg; lin2025boundary; rojas2025improving; zhu2025enhancing; wang2025revolutionizing; zhan2025principled), and have demonstrated promising results in further enhancing the reasoning capabilities of dLLMs. The key distinction from our approach is that these methods use a fixed sampling strategy (e.g., high-confidence sampling), and the policy corresponds to the dLLM itself (or a LoRA-augmented version for efficiency). Closest to our work are DCOLT (huang2025reinforcing), which trains a separate unmasking module (with a fixed number of unmasked tokens in each step) in addition to updating the base model via RL, and DiFFPO (zhao2025diffpo), which jointly learns an unmasking confidence threshold while updating the base model through RL. Differently to both these works, our primary goal is not in improving the reasoning abilities of the underlying dLLM; hence we keep the underlying dLLM fixed and focus solely on training a standalone policy with the aim of learning fast and adaptive sampling while preserving the performance of the base model. Another related concurrent work is hong2025improving where an unmasking policy for dLLMs trained via GRPO is proposed, similar to our approach. However, their policy unmasks a fixed number of tokens at each step (similar to DCOLT), whereas ours dynamically adapts the number of tokens to unmask via our Bernoulli formulation. Finally, Seed Diffusion (song2025seed) identifies computation-aware reinforcement learning as a key ingredient for achieving competitive efficiency in closed-source coding dLLMs.
@@
Besides the aforementioned heuristic-based sampling innovations, other efforts to improve inference efficiency in dLLMs (ma2025dinfer) include KV caching (jiang2025d; fastdllm2025), (variants of) speculative decoding (israel2025accelerating; campbell2025self; guo2025reviving; wu2025free), training separate decoder modules during pretraining (liu2024think; arriola2025encoder), and diffusion forcing (wang2025diffusion), among others. Concurrent work explores distilling faster generation patterns directly into the base model (chen2025dparallel), or training a (small) separate unmasking network on top of the pretrained dLLM (bansal2025enabling; bao2025learning). Crucially, unlike our approach, where this training is done via RL, these works train unmasking modules by distilling generation trajectories from the base model. As a result, we expect that such approaches may still inherit limitations observed in dLLM sampling, such as difficulties in the non semi-AR regime (cf. Section˜2.2).
@@
Since our sampling policies result in a variable number of sampling steps ($T-\hat{T}$) per input, our work connects to the broader literature on adaptive computation (graves2016adaptive; teerapittayanon2016branchynet). Notable examples of using RL to learn input-dependent compute policies include conditional computation with stochastic gating (bengio2015conditional), dynamic block skipping in residual networks (wang2018skipnet), and RL-based early-exit decisions for long chain-of-thought reasoning (dai2025s). To the best of our knowledge, our work is the first to explore learning adaptive policies using RL for the task of sampling in dLLMs. The closest related concurrent work is DiFFPO (zhao2025diffpo); however, unlike our approach–where the policy is learned end-to-end–DiFFPO learns to predict adaptive thresholds, which are then used in the same fashion as the fixed thresholds employed by Fast-dLLM (fastdllm2025). Concurrently, there is also a growing body of work that uses RL to introduce adaptivity into reasoning chains of autoregressive LLMs (arora2025training; yi2025shorterbetter).
@@ 6 Conclusion
We introduced a reinforcement learning approach for learning unmasking strategies in diffusion LLMs. Our experiments demonstrated that the learned policies can match or even exceed the performance of recently proposed sampling heuristics, paving the way for the automated discovery of scalable and robust sampling mechanisms.
Limitations and future work.
One notable limitation of our approach is that it requires training a separate policy for each $\alpha$, and that this hyper-parameter sometimes does not yield as much fine-grained control over the policy’s behavior as desired (Section˜4.1). Thus, an important future direction is to investigate whether the accuracy-speed tradeoff could be controlled through some alternative, ideally test-time, mechanism. We also observe that the optimal policy temperature $\tau_{\pi}$ varies across generation settings (Figure˜14), suggesting the need to explore whether $\tau_{\pi}$ can be learned jointly with the policy. Beyond overcoming these limitations, promising future extensions are (i) expanding the training mixture to incorporate data from multiple domains; (ii) extending our sampling policies to also support remasking (wang2025remasking; huang2025don), and (iii) moving beyond text to learn samplers for multimodal discrete diffusion (swerdlow2025unified; zhou2025co).
@@
  • In Appendix A, we provide an algorithm for our proposed policy sampling.
  • In Appendix B, we present additional plots to supplement the experiments from the main paper:
    • In B.1, we replicate Figure˜1 across models (LLaDA, Dream) and datasets (GSM8k, MATH).
    • In B.2, we replicate Figure˜4 using latency (wall-clock time) as the efficiency measure.
    • In B.3, we replicate Figure˜4 for policies trained on Dream.
    • In B.4, we show results for the same setting as 4(a) and 4(c) but with a denser $\alpha$-grid. We also show the spread of policies across different training seeds.
    • In B.5, we plot the impact of varying the policy temperature $\tau_{\pi}$.
    • In B.6, we provide model transfer results for the MATH dataset.
    • In B.7, we provide generation length (L) transfer results for the MATH dataset.
  • In Appendix C, we perform a qualitative analysis of our learned sampling policies.
  • In Appendix D, we visualize token generation trajectories.
  • In Appendix E, we describe dynamic Plackett-Luce sampling as an alternative to Bernoulli sampling.
  • In Appendix F, we detail our expert steering approach.
  • In Appendix G, we discuss an extended background on discrete diffusion models.
  • In Appendix H, we provide implementation and hyperparameters details.
  • In Appendix I, we visualize the architecture of our sampling policy.
Appendix A Policy Sampling Algorithm
Algorithm 1 Policy Sampling (full-diffusion setting $BL=L$)
1:Input: Prompt ${\bm{x}}\in{\mathcal{V}}^{d}$, maximum generation length $L$, maximum diffusion steps $T$, dLLM $p_{\theta}$, sampling policy $\pi_{\phi}$ (via $f_{\phi}$), policy temperature $\tau_{\pi}$
2:Output: Generated sequence $^{\bm{y}}_{\hat{T}}$, number of sampling steps
[…]
5:for $t=T,T-1,\ldots,1$ do
6:  if ${\mathcal{M}}_{t}=\emptyset$ then
7:   break $\triangleright$ Stop generating when all unmasked
8:  end if
9:  $p_{t}^{k}\leftarrow p_{\theta}(\cdot\mid{\bm{x}},{\bm{y}}_{t})$, $\forall k\in[L]$ $\triangleright$ Token distributions
10:  $c_{t}^{k}\leftarrow\max_{v\in{\mathcal{V}}}p_{t}^{k}(v)$, $\forall k\in[L]$ $\triangleright$ Token confidences
11:  ${\bm{c}}_{t}\leftarrow(c_{t}^{1},\ldots,c_{t}^{L})$
12:  $m_{t}^{k}\leftarrow\mathbf{1}[k\in{\mathcal{M}}_{t}]$, $\forall k\in[L]$
13:  ${\bm{b}}_{t}\leftarrow f_{\phi}({\bm{c}}_{t},{\bm{m}}_{t},t)$ $\triangleright$ Policy logits
14:  $s_{t}^{k}\leftarrow\sigma(b_{t}^{k}/\tau_{\pi})$, $\forall k\in{\mathcal{M}}_{t}$
15:  $u_{t}^{k}\sim\text{Ber}(s_{t}^{k})$, $\forall k\in{\mathcal{M}}_{t}$ $\triangleright$ Unmasking decisions
16:  $u_{t}^{k}\leftarrow 0$, $\forall k\notin{\mathcal{M}}_{t}$ $\triangleright$ Ignore already-unmasked
17:  if $\sum_{k=1}^{L}u_{t}^{k}=0$ then
18:   $u_{t}^{k^{*}}\leftarrow 1$, $k^{*}=\textrm{argmax}_{k\in{\mathcal{M}}_{t}}s_{t}^{k}$ $\triangleright$ Argmax fallback
19:  end if
20:  ${\mathcal{U}}_{t}^{\pi}\leftarrow\{k\in{\mathcal{M}}_{t}\mid u_{t}^{k}=1\}$
21:  $y_{t-1}^{k}\sim p_{t}^{k}$, $\forall k\in{\mathcal{U}}_{t}^{\pi}$ $\triangleright$ Sample tokens
22:  $y_{t-1}^{k}\leftarrow y_{t}^{k}$, $\forall k\notin{\mathcal{U}}_{t}^{\pi}$
23:  ${\mathcal{M}}_{t-1}\leftarrow{\mathcal{M}}_{t}\setminus{\mathcal{U}}_{t}^{\pi}$
24:end for
[…]
26:return $^{\bm{y}}_{\hat{T}}$, $T-\hat{T}$
Appendix B Additional Results
B.1 Figure 1 replicated for {LLaDA-8B-Instruct, Dream-7b-Instruct} $\times$ {GSM8k, MATH-500}